how to solve:7t-4(2t+3t)4/2=
Answer by blacksaibot(4)

Assuming you wrote the problem correctly... it will be the same as:
(7t-4) * (2t+3t) * (4/2)

(4/2) is easily rewritten as 2

Do what you can within the parenthesis first.
(2t + 3t) = 5t

(7t-4) * (5t) * (4/2)

Multiply the first expression (7t-4) * (5t), worry about 2 at the end

(7t * 5t) + (-4 * 5t) = (35t^2 - 20t)

So now we have: (35t^2 - 20t) * 2 = 70t^2 - 40t

Since you never wrote a number after the equal sign, you cannot really solve for t in this situation. Traditionally, you would have complete the square but again, there is nothing else to solve for.